# Data Caricatures: On the Representation of African American Language in Pretraining Corpora

### v1

@@ Appendix B Additional Data Details
Figure 6: Distribution of metadata from selected corpora.
@@ Appendix C AAL Probability Threshold
Figure 7: Full distribution of demographic alignment classifier posterior probabilities across all TwitterAAE Model Labels for C4 and all sampled corpora datacomp-lm-baseline, Redpajama-Data-V2, and fineweb, as captured by t-digest Davidson-Pilon (2015).
Any record where AAL has the greatest probability of all labels or where AAL exceeds a posterior probability of 0.3 is considered for analysis. This threshold is chosen for the individual feature analyses to capture documents that may contain features of AAL although largely be composed of WME. At the same time, this threshold does not add significant documents to the extracted dataset that would make analysis infeasible with the available computational resources. Figure 7 shows the distribution of document AAL probabilities as well as the probabilities of the other language varieties considered by the demographic alignment classifier Blodgett et al. (2016). For C4 (top), considering lower thresholds sharply increases the number of documents considered.
Appendix D Human Judgments
D.1 Annotation Details
We recruit 3 annotators that self-identify as AAL-speakers and have previous or current experience in natural language processing or linguistics. Each annotator is initially provided 420 texts total, with 50 texts shared among all annotators. Two annotators completed judgments for all 420 texts, and the remaining annotator completed 299 texts. Annotators were compensated at a rate of \$22.50 per hour.
D.2 Human Judgments Sample
Because much of recent work relies on a threshold of 0.8 to identify AAL documents (e.g., Davidson et al. (2019)) using the demographic alignment classifier Blodgett et al. (2016), we prioritize texts with a higher probability of including AAL in human judgments. Additionally, because C4.en.noBlocklist is not used for pretraining, we prioritize texts from the filtered subset of C4 (C4.en). The composition enforced on the sampled human judgment texts is shown in Table 6. Notably, only 10% of the texts in the sample have $\geq 90\%$ probability because there are few texts in the full extracted dataset that meet this requirement. Considering this sampling, all reported estimates based on human judgments (e.g., overall Stereotype estimates) are calculated through a weighted average across probability ranges, weighted by the size of each probability range in the full corpus.
AA Prob % Sample
.$.9\leq P_{\_}{AA}.$ 0.1
.$9\displaystyle.8\geq P_{\_}{AA}\geq.$9 0.35
.$8\displaystyle.7\geq P_{\_}{AA}\geq.$8 0.3
.$7\displaystyle.6\geq P_{\_}{AA}\geq.$7 0.15
.$6\displaystyle.5\geq P_{\_}{AA}\geq.$6 0.1
@@ D.3 Annotator Agreement
Dimension $\kappa$ $\kappa_{\_}{bin}$ $r_{\_}s$ Support
Human 0.226 0.581 0.588 40.3
@@
Table 7: Average pairwise interannotator agreement results for each dimension using Cohen’s $\kappa$ and Spearman $r_{\_}s$. Cohen’s $\kappa_{\_}{bin}$ represents agreement with binarized judgments ($\{0,1\}\rightarrow 0$, $\{2,3\}\rightarrow 1$).
Table 7 presents the inter-annotator agreement scores for each judgment dimension. Annotators show moderate to substantial agreement for binarized Humanness ($\kappa_{\_}{bin}=0.581$), Dialect ($\kappa_{\_}{bin}=0.747$), and Native Speaker ($\kappa_{\_}{bin}=0.619$), as well as strong Spearman correlations.
@@ D.4 Annotation Interface
Because we are interested in capturing the annotators’ own perspectives on the pretraining texts and their interpretations of these dimensions, we avoid providing detailed definitions of the dimensions–particularly the Stereotype and Appropriateness dimensions. We also include an additional space for comments, including reasoning for selected judgments as well as if the annotator recognizes a given text from a different source. Screenshots of the interface are provided in Figure 8, Figure 9, and Figure 10.
@@ Appendix E Feature Abbreviations
Appendix F Hip Hop n-gram Analysis
@@
Appendix G Additional Hip Hop Analysis
@@ Appendix H Non-AAL Annotations
Appendix I Automated Filters
I.1 Filter Details
@@ I.2 Replicated Filters
PALM Quality Filter. The PALM quality filter Chowdhery et al. (2022) uses Wikipedia samples, OpenWebText, and RedPajama-v1 books as the positive class and Common Crawl as the negative class. We randomly sample Because the books subset of RedPajama-v1 is unavailable, we instead use the Project Gutenberg books corpus.  [^11] https://huggingface.co/datasets/manu/projectgutenberg We use publicly available corpora for Wikipedia  [^12] https://huggingface.co/datasets/wikimedia/wikipedia and OpenWebText  [^13] https://huggingface.co/datasets/Skylion007/openwebtext Gokaslan et al. (2019). We sample 600 CCNet snapshots weighted among head (10%), middle (20%), and tail (70%) partitions made available through and following RedPajama-v2 Weber et al. (2024). We sample 250,000 positive and negative references, with positive references split evenly among the three corpora.
@@
Appendix J AAL Sources in Pretraining Corpora

### v2

@@ Appendix B Additional Data Details
Figure 7: Distribution of row count, compression ratio, file size, and bytes per record from select corpora. This captures the variety of metadata standards between the different corpora.
@@ Appendix C AAL Probability Threshold
Figure 8: Full distribution of demographic alignment classifier posterior probabilities across all TwitterAAE Model Labels for C4 and all sampled corpora DCLM-baseline, RedPajama-Data-V2, and FineWeb, as captured by t-digest Davidson-Pilon (2015).
Any record where AAL has the greatest probability of all labels or where AAL exceeds a posterior probability of 0.3 is considered for analysis. This threshold is chosen for the individual feature analyses to capture documents that may contain features of AAL although largely be composed of WME. At the same time, this threshold ensures that the extracted dataset is feasible to analyze with the available computational resources. Figure 8 shows the distribution of document AAL probabilities as well as the probabilities of the other language varieties considered by the demographic alignment classifier Blodgett et al. (2016). For C4 (top), considering lower thresholds sharply increases the number of documents considered.
Appendix D Human Judgments
D.1 Annotation Details
We recruit 3 annotators that self-identify as AAL speakers and have previous or current experience in natural language processing or linguistics. Each annotator is initially provided 420 texts total, with 50 texts shared among all annotators. Two annotators completed judgments for all 420 texts, and the remaining annotator completed 299 texts. Annotators were compensated at a rate of \$22.50 per hour.
D.2 Human Judgments Sample
[…]
AA Prob % Sample
@@ D.2 Human Judgments Sample
[…]
@@ D.3 Annotator Agreement
[…]
Human 0.226 0.581 0.588 121
@@
[…]
@@ D.4 Annotation Interface
We also include an additional space for comments, including reasoning for selected judgments as well as if the annotator recognizes a given text from a different source (e.g., song lyrics). Screenshots of the interface are provided in Figure 9, Figure 10, and Figure 11.
@@ Appendix E Feature Abbreviations
Appendix F AAL Classifier Examples
Corpus Text AAL Prob.
Dolmino (Dolmino-mix) …Title: heal yea Review: i b chillin wit da dawg 4 real luda iz da tytest rhyma in da hystory of da word 4 eva!! He got mad skillz cuz he keep rhymin bout hiz hoez which iz tyte cuz my whyte azz dawgs keep chillin at my plaze wit dis CD YO!! It real tyte how he spell mouf 4 real… 0.91
RedPajama-v1 …POPULAR: wiz khalifa, moneybagg yo, young jeezy, kevin gates, lil wayne, curren\$y, fabolous, nba youngboy, eminem, mac miller, meek mill, lil baby, future 0.80
[…]
FineWeb-Edu …"Duh wite root," pointing to a wild shrub, "dey use fuh stomach troubles. Buttuh root an palmettuh root an May apple, yuh bile tuhgedduh wid a quawt uh watuh till it simmuh down tuh haf uh pint, den yuh add some cawn wisky. Dat a fambly tonic tuh buil yuh up."… 0.51
Table 9: Abbreviated examples of texts and accompanying posterior probabilities of AAL according to the demographic alignment classifier. Documents are sampled from each of 5 bins between 0.5 and 1.0 including all corpora. Texts are shown exactly as they appear in corpus documents.
Corpus Text AAL Prob.
[…]
Table 10: Abbreviated examples of texts and accompanying posterior probabilities of AAL according to the demographic alignment classifier. Documents are sampled from each of 5 bins between 0.5 and 1.0 including all corpora.
Table 9 presents randomly sampled abbreviated texts from the AAL subsets of the corpora analyzed. As shown in Table 2, the proportion of documents determined to contain identifiable AAL features by annotators is low for high posterior probability ranges. Considering the demographic alignment classifier Blodgett et al. (2016) uses token-level features, some of these high AAL probability documents contain repeated abbreviations while others contain dense usage of AAL as shown in the top row of Table 9.
Appendix G Full Feature Analyses
We measure the distribution of AAL features across all corpora using random samples of 50,000 documents  [^10] Some AAL subsets of corpora are smaller than 50,000. For these corpora, we consider all documents. meeting the AAL threshold probability of 0.3. While there are some commonalities among corpora, such as Zero Plural (ZP) and Zero Copula (ZC) appearing frequently related to most other features, the frequency of individual features and order of most frequent features varies widely (e.g., Remote done (RD) is highly frequent in OpenWebText, but not in others.)
Figure 12: Feature distribution over documents for all corpora. Shaded grey region indicates the average frequency across all corpora.
Appendix H Hip Hop n-gram Analysis
@@
Appendix I All Corpora Hip Hop Analysis
We analyze a 50% random sample [^12] We consider all of the AAL texts extracted from OpenWebText given its size. of all pretraining corpora and measure the presence of n-grams shared with hip hop lyrics. Figure 13 presents the results of the hip hop lyrics analysis for all corpora. The proportion of documents containing hip hop n-grams varies widely across corpora, with 13-gram hip hop sequences appearing in between roughly 2-50% of documents. Overall, however, we see that language resembling hip hop is substantially represented in each corpus.
Figure 13: % of AAL documents extracted from each corpus containing shared n-grams with hip hop lyrics for varying n-gram token lengths. Tokenization uses the Llama-2 tokenizer.
Appendix J Additional Hip Hop Analysis
@@ Appendix I All Corpora Hip Hop Analysis
Appendix L Automated Filters
L.1 Filter Details
@@
PALM Quality Filter. The PALM quality filter Chowdhery et al. (2022) uses Wikipedia samples, OpenWebText, and RedPajama-v1 books as the positive class and Common Crawl as the negative class. Because the books subset of RedPajama-v1 is unavailable, we instead use the Project Gutenberg books corpus.  [^14] https://huggingface.co/datasets/manu/projectgutenberg We use publicly available corpora for Wikipedia  [^15] https://huggingface.co/datasets/wikimedia/wikipedia and OpenWebText  [^16] https://huggingface.co/datasets/Skylion007/openwebtext Gokaslan et al. (2019). We sample 600 CCNet snapshots weighted among head (10%), middle (20%), and tail (70%) partitions made available through and following RedPajama-v2 Weber et al. (2024). We sample 250,000 positive and negative references, with positive references split evenly among the three corpora.
@@
Appendix M AAL Sources in Pretraining Corpora